Calculate the size of the stellar aberration effect in Newtonian physics for
light coming straight down at speed c and a telescope moving straight
sideways at speed v. At what angle θ must we point the telescope to see
the star? Hint: In what direction is the light moving in the reference frame
of the telescope? Express θ in terms of v and c. Give the exact value, not
the small v approximation mentioned in the main chapter.
1
Expert's answer
2019-10-14T09:41:22-0400

In this problem the velocity of telescope is perpendicular to that of light so,

Abberation angle will be,

"\tan\theta=\frac{v}{c}"

If v is the tangential velocity of earth around sun

And also c is speed of light

So,

"v=3\times 10^4\ m\ sec^{-1}"

"c=3\times10^8\ m\ sec^{-1}"

So,

"\tan\theta=10^{-4}"


Now

"\theta =tan\theta= 0.0001\ radian"